# Superfast Low Rank Approximation

### v2

@@ Abstract
Low rank approximation of a matrix (hereafter LRA) is a highly important area of Numerical Linear and Multilinear Algebra and Data Mining and Analysis. One can operate with an LRA superfast – by using much fewer memory cells and flops than an input matrix $M$ has entries.111“Flop” stands for “floating point operation”. Can we, however, compute an LRA of a matrix superfast? Yes and no. For worst case inputs, any LRA algorithm fails miserably unless it involves all entries of $M$, but in computational practice worst case inputs seem to appear rarely, and accurate LRA are routinely obtained superfast for large and important classes of matrices. We test numerically a number of superfast algorithms for various synthetic and real world matrices and consistently arrive at quite accurate LRA. Adequate formal support for these empirical observations is still a challenge, but our upper estimates for the output accuracy of our superfast deterministic LRA algorithms show that they only fail for a very narrow input class and moreover that the spectral and Frobenius error norms of the outputs of our randomized LRA algorithms are expected to stay within a noticeable but reasonable factor from their optima.
@@
Low-Rank Approximation (LRA), CUR LRA, Sublinear algorithms.
@@ 1 Introduction
1.1. Superfast LRA CUR LRA: background. LRA of a matrix is among the most fundamental problems of Numerical Linear and Multilinear Algebra and Data Mining and Analysis, with applications ranging from machine learning theory and neural networks to term document data and DNA SNP data (see surveys [39, 44, 46, 69]). For example, matrices that represent Big Data (e.g., unfolding matrices of multidimensional tensors) tend to be so immense that only a tiny fraction of the entries fits primary memory of a computer, but quite typically they admit LRA, that is, can be closely approximated with the products of a few matrices of smaller sizes [72]. Such products can be multiplied by a vector superfast, that is, at sublinear cost – by using much fewer scalars and flops than the matrix has entries. [^2] Superfast algorithms have been studied extensively for Toeplitz, Hankel, Vandermonde, Cauchy, and other structured matrices having small displacement rank (see more in Sec. 1.5), and we extend this concept to computations of LRA of matrices that admit LRA. This implies dramatic acceleration of various popular matrix computations reduced to that basic operation, but can we compute an LRA of a matrix superfast or at least quasi superfast, that is, by involving only a small fraction of all entries of an input matrix?
@@
Yes and No. No, because on the worst case inputs and even on the small matrix families of Example A.1 any algorithm fails miserably for computing LRA and even for a posteriori estimation of its accuracy unless it involves all entries of an input matrix. Yes, because in computational practice worst case inputs seem to appear rarely, and accurate LRA are routinely computed superfast for a large and important class of matrices. Most popular among these algorithms are deterministic Cross-Approximation (C–A) iterations, specializing Alternating Directions Implicit (ADI) Method [63] to LRA and recalled in Sec. 3.2. They output LRA in a memory efficient form of CUR, which preserves sparsity and non-negativity of an input matrix and is widely applied in data analysis.
@@
Given an $m\times n$ matrix $M$ and a target rank 333“In practice, the target rank is rarely known in advance. $\dots$ LRA algorithms are usually implemented in an adaptive fashion $\dots$ until the error norm satisfies the desired tolerance” [70, Sec. 2.4]. $r$, one can fix a pair of small integers $k$ and $l$,
$r\leq k\leq m,~r\leq l\leq n,$ (1.1)
and then fix or randomly sample a pair of sets of $l$ columns and $k$ rows of $M$. The submatrices made up of these columns and rows serve as the factors $C$ and $R$ of a CUR of $M$, respectively, and share a $k\times l$ submatrix $G$ of $M$, said to be the generator of CUR. Given a generator $G$, one can compute a factor $U$ by using about $kl$ scalars and $O(kl\min\{k,l\})$ flops (see Sec. 2), that is, quasi superfast where $kl\ll mn$ and superfast where $kl\min\{k,l\}\ll mn$.
Our formal and empirical study of superfast LRA includes some novel algorithms and techniques. [^4] Our link of maxvol CUR rank-1 algorithm to matrix norm estimation in Sec. 3.3 can be of independent interest.
1.2. Our contribution I: the study of CUR LRA. In Part I we seek CUR LRA deterministically. In Table 5.1 we display the results of our numerical test of C-A iterations and some other superfast algorithms for computing CUR for various $n\times n$ synthetic and real world matrices for $n=256,512,1024$. We always had spectral output error norm within a factor of 30 from optimal with a single exception for the Primitive algorithm that uniformly at random sampled $k$ rows and $l$ columns of $M$ – to define factors $C$ and $R$ and a generator $G$ of CUR.555 Instead of sampling one can fix the rows and columns of $M$, e.g., the first ones, the last ones, or equally spaced ones among all $m$ rows and $n$ columns. Its output error norm has exceeded the optimal one by factors ranging from 1,000 to 40,000.
Adequate formal support for these empirical observations is a challenge, but we partly met in Thm. 4.1 (our Main Result I). According to this theorem any CUR LRA is within a factor of $(v+1)^{2}$ from optimal under the spectral norm $||\cdot||$ for
@@
For comparison, Example A.1 shows that one cannot estimate MCUR||M-CUR|| superfast directly for a worst case matrix $M$ and a fixed $CUR$. Our upper bound of order of $(v+1)^{2}$ is too high to guide practical LRA but should be of qualitative interest. Our Tables 5.2 and 5.3 show empirically that superfast combination of the randomized near-optimal algorithms of [20] and C-A iterations tends to keep the output CUR LRA nearly as accurate as in the case of the randomized near-optimal algorithms of [20], running at superlinear cost.
@@
1.3. Our contribution II: dual sketching algorithms. We provide another insight into the problem by deducing a little stronger error bound for a large class of superfast randomized LRA algorithms. With a high probability (whp) random sketching LRA algorithms [44, 39, 73, 70, 46, 69] compute near-optimal rank-$r$ approximation [^6] Here and hereafter “rank-$r$ approximation” means “approximation of rank at most rr”. of an $m\times n$ matrix $M$ by using auxiliary matrices (sketches) $FM$ and/or $MH$ for $k\times m$ and/or $n\times l$ random sketch matrices $F$ and/or $H$, respectively, for $r\leq\min\{k,l\}$, $\max\{k,l\}=r+p$, and an oversampling (overestimation) parameter $p>0$. For small positive integers $k$ and $l$ the algorithms run superfast, except for the computation of the sketches, and we seek superfast algorithms at that stage as well.
@@
To simplify our presentation, we restrict it to real matrices throughout, with $M^{*}$ denoting transpose rather than Hermitian transpose of $M$, and to random sketching LRA that only uses sketches $MH$, as in [39], although in Sec. 7.2 we extend our results to algorithms using both sketches $FM$ and $MH$, and we can readily extend our study to complex matrices.
@@
One can explore a variety of ways towards our goal. Maybe the simplest way is to randomly sample $l$ columns of the matrix $M$ or equivalently to let $H$ be the $n\times l$ leftmost submatrix of a random $n\times n$ permutation matrix under a fixed randomization model. Clearly, in this case only $nl$ elements of $M$ are accessed, and the calculation itself is given ”for free”. We can alternatively compute sketches $MH$ superfast by choosing Ultrasparse sketch matrices $H$. Example A.1 tells us that these ways (as well as any other superfast way) cannot yield accurate LRA for all inputs $M$, but in our numerical experiments with Ultrasparse sketch matrices we have quite consistently output accurate LRA for a large input class.
@@
Towards explaining these test results and constructing a theory, we introduce a probabilistic structure in the space of input matrices $M$, thus defining dual sketching in comparison to customary primal one, where this matrix is fixed.
The output error norm of sketching algorithms of [39] is essentially defined by the product $V_{1}^{*}H$ where $V_{1}$ is the matrix of the $r$ top left singular vectors of $M$, associated with its $r$ top (largest) singular values. Now assume that $V_{1}^{*}$ is the Q factor in QR or QRP factorization of a Gaussian random matrix, filled with independent identically distributed normal (standard Gaussian) random variables, $H$ has orthonormal columns, and lrl-r is reasonably large. Then we readily prove that whp the sketching algorithm outputs LRA whose both Frobenius and spectral error norms are within a factor of $\sqrt{1+16~n/l}$ from optima; this is essentially Thm. 8.2 – our Main Result 2!
Thm. 8.2 can be applied to approximation by matrices of any fixed rank $r$ and cover random sketching algorithms that in the case of a proper Ultrasparse sketch matrix $H$ use $O((m+n)r^{2})$ flops, running superfast where $r^{2}\ll\max\{m,n\}$ and quasi superfast already where $r\ll\min\{m,n\}$.
1.4. Quality of dual LRA. Like Thm. 4.1, Thm. 8.2 should be of some qualitative interest, e.g., it implies that miserable failure of our superfast dual sketching algorithms is highly unlikely under our probabilistic model, although the algorithms fail for worst case inputs.
On the other hand, the upper bound of Thm. 8.2 is still too high, and we propose and successfully test some heuristic improvement of output accuracy (see Sec. 9):
(i) Thm. 8.2 holds for any matrix $H$ that has orthonormal columns or is just well-conditioned. To increase our chances for obtaining accurate LRA we apply sketching algorithms with a number of distinct sketch matrices $H$ or their combinations.
@@
(ii) For sketch matrices obtained with sparse subspace embedding [9, 12, 14], [71, Sec. 3.3], [46, Sec. 9], one can devise LRA algorithms running at linear computational cost. According to [40], such acceleration versus LRA running at superlinear cost with Gaussian sketch matrices tends to make the output accuracy of the LRA somewhat less reliable, although [14] partly overcomes this problem for incoherent matrices, that is, filled with entries of comparable magnitude (see [14] or [11, Def. 1.1] for formal definition of incoherence). Multiplication by the matrix of Subsampled Randomized Hadamard or Fourier Transform (SRHT or SRFT) tends to make a matrix $M$ incoherent [14] but at superlinear cost. For a compromise, we devise Abridged SRHT sketch matrices. They are Ultrasparse and can be multiplied by a dense matrix superfast. Furthermore, they have orthonormal columns, and so Thm. 8.2 provides some probabilistic motivation for this otherwise heuristic LRA; our tests provide its empirical support. [^7] Abridged SRFT matrices can be defined similarly to Abridged SRHT matrices and empirically have similar power (see [55, Sec. 13.6 and Appendix B]); unlike real Abridged SRHT matrices they involve nonreal roots of unity.
1.5. Related work. For two decades of their worldwide application C-A (ADI-like) iterations have been consistently computing accurate CUR LRA superfast for a large and important matrix class (see [66, 36, 37, 34, 67, 2, 32, 6, 5, 30, 3, 33, 49, 1], and the bibliography therein). More recently, randomized superfast LRA algorithms have been proposed in [47, 8, 13] for some specified large and important classes of matrices defined by large numbers of independent parameters. Most notably, the authors of [47, 13] proved that their algorithms are expected to output near-optimal LRAs for Symmetric Positive Semidefinite (SPSD) matrices.
The superfast algorithms of [47, 8, 13] output LRA but not CUR LRA. The known algorithms extend any LRA to CUR LRA – at the price of increasing the spectral norm bounds on the output errors by a factor of $\sqrt{1+(n-r)r)}$ (see [39, Secs. 3.3.2 and 3.3.3]).
@@
The sublinear cost of the algorithms of [47, 8, 13] does not include the superlinear cost of a posteriori estimation of their output error norms and correctness verification. The superfast deterministic algorithm of [42, Part III], however, computes CUR LRA of an $n\times n$ SPSD matrix with both spectral and Frobenius error norms within a factor of $n$ from optimal and as by-product, at no additional cost, estimates the output error norm, verifying correctness.
[43, Alg. 2.2] is a heuristic superfast CUR LRA algorithm, which is accurate for a large class of unsymmetric matrices according to the test results in [43]. The algorithm combines uniform choice of some subsets of the column and row sets for CUR with their updating by means of the Interpolative Decomposition of [39, Sec. 3.2.3] based on QR factorization with column pivoting. Jianlin Xia in [74] modified this algorithm by using Strong Rank Revealing (SRR) QR factorization of [26] instead of QR factorization with pivoting, called this technique “progressive alternating direction pivoting”, combined it with his additional techniques, in particular, randomized error estimation, and consistently obtained highly accurate LRAs for various real world inputs in his extensive numerical tests. Xia cited [42, 58] as his close predecessors, but our present algorithms rely on distinct techniques.
In formal study of CUR LRA, quite surprising was the theorem by Zamarashkin and Osinsky in [76] that for any $m\times n$ matrix $M$, any positive integer $k\leq\min\{m,n\}$, there exists a $k\times k$ CUR generator $G$ that defines CUR, which is optimal up to a factor of $k+1$ approximation of $M$ under the Frobenius matrix norm. In the next major step, Alice Cortinovis and Daniel Kressner [15] computed such a CUR at deterministic cost $O(kmn^{3})$ for $m\geq n$ by linking the problem to intricate subset selection. Superfast computation of a good generator, however, is a challenge, and the error norms of CUR LRAs of $M$ are unbounded over ill-conditioned generators $G$.
Chiu and Demanet in [11] proved that under the uniform sampling of row and column sets of $M$ the Primitive algorithm outputs accurate CUR LRA of $M$ whp, provided that there exists an accurate LRA $XYZ\approx M$, e.g., defined by the SVD of $M$, such that the matrices $X$ and $Z^{*}$ have orthonormal columns and are incoherent. [11] also extended these estimates to the case where only a single factor $X$ in some LRA $XYZ\approx M$ has orthonormal columns and is incoherent provided that the Primitive algorithm is extended with a single C-A iteration.
Alice Cortinovis and Lexing Ying [19] deduce the results of [11] under weaker assumptions about the matrices $X$ and $Z^{*}$ – they allow some of the column vectors to be sparse rather than incoherent if CUR LRA algorithm incorporates progressive alternating direction pivoting of [43, 74].
@@
The highly important estimates of [11, 19] are still limited by the basic provision of incoherence of the factors of some unknown LRA of $M$. This provision does not hold for a large class of inputs and generally cannot be verified superfast for a fixed $M$ – in contrast to the bound of Thm. 4.1.
In Parts I and II we extend the unpublished preprints [54, 55, 56, 57], which cited [59, 60, 61] as their predecessors and stated motivation for further study of superfast LRA as their major goal; this is a major goal of our current work as well. Encouraged by the success of [74, 19] we revisit the two preprints [56, 57] and extract from them our Main Results I and II and the test results of Parts I and II, respectively.

### v3

# Novel Insights into Low Rank Approximation at Sublinear Cost

@@ Abstract
Low rank approximation of a matrix (LRA) is a highly important area of Numerical Linear and Multilinear Algebra and Data Mining and Analysis. One can operate with an LRA superfast – by using much fewer memory cells and flops than an input matrix has entries.111“Flop” stands for “floating point operation”. Can we, however, compute an LRA of a matrix superfast? YES and NO. For worst case inputs, any LRA algorithm fails miserably unless it involves all input entries, but in computational practice worst case inputs seem to appear rarely, and accurate LRA are routinely computed superfast for large and important classes of matrices, in particular in the memory efficient form of CUR, widely used in data analysis. We advance formal study of this YES and NO coexistence by proving novel universal upper bounds on the spectral output error norms of all CUR LRA algorithms and, under a fixed probabilistic structure in the space of input matrices, on both spectral and Frobenius error norms of nearly all sketching LRA algorithms. These bounds imply that superfast LRA algorithms of the two kinds fail miserably only for a very narrow input class. Furthermore, in our numerical tests such superfast algorithms were consistently much more accurate than our upper estimates ensure and usually were reasonably close to optimal.
@@
Low-Rank Approximation, CUR LRA, Sublinear algorithms.
@@ 1 Introduction
1.1. Superfast LRA CUR LRA: background. LRA of a matrix is among the most fundamental problems of Numerical Linear and Multilinear Algebra and Data Mining and Analysis, with applications ranging from machine learning theory and neural networks to term document data and DNA SNP data (see surveys [30, 37, 59]). For example, matrices that represent Big Data (e.g., unfolding matrices of multidimensional tensors) tend to be so immense that only a tiny fraction of their entries fits primary memory of a computer, but quite typically they admit LRA, that is, can be closely approximated with the products of a few matrices of smaller sizes [62]. Such products can be multiplied by a vector superfast, aka at sublinear cost, – by using much fewer memory cells and flops than the matrix has entries. [^2] Superfast algorithms have been studied long and extensively for structured matrices having small displacement rank (see more in Sec. 1.4.1), and we extend this concept to computations of LRA. This implies dramatic acceleration of various popular matrix computations reduced to that basic operation, but can we compute an LRA of a matrix superfast or at least quasi superfast, that is, by involving only a small fraction of all entries of an input matrix? YES and NO.
NO, because on the worst case inputs and even on the small matrix families of our Example A.1 any algorithm fails miserably for computing accurate LRA and even for a posteriori estimation of its accuracy unless it involves all entries of an input matrix.
@@
YES, because in computational practice worst case inputs seem to appear rarely, and accurate LRAs are routinely computed superfast for a large and important class of matrices, in particular by means of the popular Cross-Approximation (C–A) iterations, specializing the Alternating Directions Implicit (ADI) method [54] to LRA and recalled in Sec. 3.2.
Their output LRAs are memory efficient products CUR where the factors $C$ and $R$ are made up of $l$ columns and $k$ rows of an input matrix, respectively, the factor $U$ has size $k\times l$,
$r\leq k\leq m,~r\leq l\leq n,~kl\ll mn,$ (1.1)
@@
and $r$ denotes a target rank,333“In practice, the target rank is rarely known in advance. $\dots$ LRA algorithms are usually implemented in an adaptive fashion” – with submatrices $R$ and $C$ recursively expanded “until the error norm satisfies the desired tolerance” [60, Sec. 2.4]. presumed to be low. The $k\times l$ submatix $G$ shared by the factors $C$ and $R$ is said to be CUR generator. Given $M$ and $G$, the factors $C$ and $R$ are immediately available, and one obtains canonical CUR by computing the factor $U$ in about $kl$ scalars and $O(kl\min\{k,l\})$ flops (see (2.3)). This computation is quasi superfast for $kl\ll mn$ and superfast for $kl\min\{k,l\}\ll mn$.
In [67] Zamarashkin and Osinsky proved that, for any $m\times n$ matrix $M$ and any positive integer $k\leq\min\{m,n\}$, there exists a CUR LRA defined by a $k\times k$ generator and optimal up to a factor of $k+1$ under the Frobenius matrix norm. In [13] Cortinovis and Kressner computed such a CUR at deterministic cost $O(kmn^{3})$ for $m\geq n$ by means of linking the task to subset selection. Superfast computation of a good generator, however, is a challenge, and the error norms of CUR LRAs of $M$ are unbounded for ill-conditioned generators $G$.
1.2. Our contribution. We extract our Part I and Part II from preprints [47, 48], respectively (see further comments on this in Sec. 1.4.3), and point out applications of LRA to superfast matrix norm estimation and least squares approximation of multivariate functions in Secs. 3.3 and 11.
1.2.1. Part I: the study of CUR LRA. According to our Main Result 1 (Thm. 4.1), any CUR LRA (no matter how computed) has spectral output error norm $||\cdot||$ within a factor of $(v+1)^{2}$ from optimal for
@@
This universal upper bound implies that any CUR LRA can fail miserably only on a narrow class of inputs. In our tests various known and novel superfast CUR LRA algorithms were consistently much more accurate than this upper bound ensures and usually have spectral output error norms within a factor of 30 from optima for $n\times n$ inputs and $n=256,512,1024$ (see Table 6.1).
Furthermore, instead of customary combinations of C-A iterations Strong Rank Revealing (SRR) QR or LU factorization of [19, 41], we tested their combination with random sampling algorithms of [16], and in our tests the output accuracy of this superfast CUR LRA was almost always nearly as accurate as in the case of the algorithms of [16] themselves, which run at superlinear cost and are expected to be near-optimal under the Forbenius norm (see Tables 6.2 and 6.3).
@@
1.2.2. Part II: sketching LRA algorithms. We provide another insight into this subject by proving a little stronger error bound for a large class of superfast LRA algorithms. First recall that with a high probability (whp) random sketching LRA algorithms of [64, 60] compute near-optimal rank-$r$ approximation [^4] Here and hereafter “rank-$r$ approximation” means “approximation of rank at most rr”. of an $m\times n$ matrix $M$ under the Frobenius norm by using auxiliary matrices (sketches) $FM$ and/or $MH$ for proper $k\times m$ and/or $n\times l$ random multipliers $F$ and/or $H$, respectively, $\max\{k,l\}=r+p$, $k,l,m,n,r$ satisfying (1.1), and a positive oversampling (overestimation) parameter $p\ll\min\{m,n\}$. Usually these algorithms output LRA but not CUR LRA (see [16] for an important exception), but one can extend any LRA to CUR LRA – at the price of increasing the spectral norm bounds on the output errors by a factor of $\sqrt{1+(n-r)r)}$ (see [30, Secs. 3.3.2 and 3.3.3]).
Apart from the stage of the computation of sketches, random sketching algorithms of [30] (resp. of [60]) and their extensions in our Sec. 7 run superfast where $l^{2}\ll n$ (resp. $l^{2}\ll n$ and $k^{2}\ll m$) and quasi superfast where $l\ll n$ (resp. $l\ll n$ and $k\ll m$). We compute the sketches superfast as well by choosing Ultrasparse multipliers $F$ and/or $H$ (see Appendix B.2). This way formal support for near-optimality of LRA whp is lost, but next we specify our weaker but meaningful alternative.
@@
To simplify our presentation, we restrict it to real matrices throughout, with $M^{*}$ denoting transpose rather than Hermitian transpose of $M$, and to the random sketching LRA that only uses sketches $MH$, as in [30], while in Sec. 8.2 we extend our results to algorithms that use both sketches $FM$ and $MH$.
@@
One can explore a variety of ways towards our goal. Maybe the simplest way is to randomly sample $l$ columns of the matrix $M$ or equivalently to let $H$ be the $n\times l$ leftmost submatrix of a random $n\times n$ permutation matrix under a fixed randomization model; in this case only $nl$ elements of $M$ are involved, and the calculation itself is given ”for free”. We, however, consider the more general class of all sketches $MH$ that can be computed superfast. Example A.1 tells us that we cannot yield accurate LRA for worst case inputs $M$ superfast, but in our numerical experiments with Ultrasparse multipliers $F$ and $H$ we have consistently output quite accurate LRA for a large class of matrices $M$.
@@
Towards explaining these test results and constructing a theory, we introduce a probabilistic structure in the space of input matrices $M$ rather than in the space of multipliers $H$, thus defining dual sketching in comparison to customary primal one, where matrix $M$ is fixed and multiplier $H$ is random.
[…]
In Sec. 9.3 we study dual LRAs under another choice of probabilistic structure, which may be more relevant to LRA computation in the real world: we fix a multiplier $H$ with orthonormal columns and let $M:=A\Sigma B+E$ for a perturbation matrix $E$ and for $A\Sigma B$ being a random pseudo SVD of a rank-$r$ matrix of size $m\times n$, said to be a two-sided factor-Gaussian (with expected rank $r$). Namely, we let $\Sigma$ be any $r\times r$ matrix of full rank $r$ (e.g., a diagonal matrix with $r$ positive diagonal entries, as in SVD) and let $A$ and $B$ be scaled Gaussian matrices, rather than matrices of singular vectors of SVD. [^5] By saying “LRA” we assume that $r$ is much smaller than $\min\{m,n\}$; then we can motivate the above definition of random pseudo SVD by recalling (e.g., from [18, Theorem 7.3] or [53]) that $\frac{1}{\sqrt{k}}G$ is close to a matrix with orthonormal columns whp for $r\ll k$ and an $n\times r$ Gaussian matrix $G$. Most of our study (in particular Theorem 9.1 and our Main Result 3, outlined below) applies to a more general class of perturbed right factor-Gaussian matrices, of the form $AG+E$ where $G$ is a $r\times n$ Gaussian matrix and $A$ is an $m\times r$ matrix of full rank $r$; transposition extends the study to perturbed left factor-Gaussian matrices.
For a matrix $M$ with a certain gap in the spectrum of its singular values [^6] This assumption is restrictive but does not exclude highly important kernel matrices [11, 65]. we prove that our superfast sketching algorithms output close rank-$r$ approximations whp on the defined probability space: namely, our bound on the output errors under this randomization model only increases from a factor of $\sqrt{1+16~n/l}$ versus the optimum to a factor of $\sqrt{1+100~n/l}$, and this is essentially Thm. 9.2 – our Main Result 3. Over the same class of factor-Gaussian matrices we also strengthen the bound of our Main Result 1 whp (see Remark 5.1).
In our tests we sampled random multipliers $F$ and $H$, even though our Main Results 2 and 3 (Thms. 9.1 and 9.2) also hold where these multipliers are fixed.
@@
1.3. Quality of superfast dual LRA. Can we decrease the upper bounds of Thms. 9.1 and 9.2 if we choose multipliers $H$ according to a proper heuristic policy? Such a policy can be shaped empirically, based on testing various candidate multipliers $H$, but here is our semi-heuristic considerations. The best-studied LRA algorithms use Gaussian multipliers $F$ and $H$, output near-optimal LRA whp, but run at superlinear cost. With sketches obtained by using sparse subspace embedding [10, 12], [61, Sec. 3.3], [37, Sec. 9], one devises LRA algorithms running at linear computational cost. According to [32], such acceleration tends to make the output accuracy somewhat less reliable. This problem is partly overcome for incoherent matrices, whose all entries have comparable magnitude (see [12] for this study and formal definition of incoherence). Multiplication by the matrix of Subsampled Randomized Hadamard or Fourier Transform (SRHT or SRFT) tends to make a matrix $M$ incoherent [12] but at superlinear cost.
Now, for our heuristic recipe, we devise $d$-Abridged SRHT multipliers in Appendix B.2.777Abridged SRFT matrices can be defined similarly to Abridged SRHT matrices and empirically have similar power (see [46, Sec. 13.6 and Appendix B]); unlike real Abridged SRHT matrices they involve nonreal roots of unity. For $d=\log_{2}(n)$ they turn into $n\times l$ SRHT multipliers, but for small positive integers $d$ (we use $d=3$ in our tests) they are Ultrasparse and can be multiplied by a dense matrix superfast. Furthermore, they have orthonormal columns, and so Thm. 9.1 provides dual probabilistic motivation for this otherwise heuristic choice of multipliers. In Sec. 10 we provide its empirical support by testing our superfast extensions of the algorithms of [30, 60]. Empirical benefits of $d$-Abridged SRHT multipliers were observed in [48] and apparently are still a novelty. In extensive numerical tests in [21] for $d=3$ they were at least as efficient as Gaussian most of the time although caused blow up of the output errors of LRA for a narrow class of inputs.
With this reservation we can use $d$-abridged SRHT multipliers instead of Gaussian towards additional support of our dual model in Appendix D. There we prove that pre- and/or post- multiplication by Gaussian matrices turns any matrix admitting LRA into a small-norm perturbation of a factor-Gaussian matrix. This multiplication becomes superfast if we substitute $3$-abridged SRHT multipliers for Gaussian.
At the end of Sec. 1.5 we comment on goals and Abridged SRHT by-product of our numerical tests.
1.4. Related work. We recall two groups of the known studies of superfast LRA.
1.4.1. Formally supported superfast LRA. Superfast algorithms have been long studied for data sparse matrices – defined by small numbers of parameters such as ones with displacement structure (cf. [31, 5, 42, 66, 43] and the bibliography therein), but some of more recent superfast algorithms output LRAs of the families of data dense matrix families, defined by order of $mn$ parameters in the space of $m\times n$ matrices. In this subsection such families are specified formally, although in some cases in terms of parameters not available superfast.
The randomized superfast algorithms of [38, 11] and [7] compute LRA for $n\times n$ Symmetric Positive Semidefinite (SPSD) and distance matrices, respectively. The superfast algorithms of [38, 7] are near-optimal whp under the Frobenius norm, while the algorithm of [11], based on random pivoted Cholesky factorization, outputs oversampled LRA of size $k\times k$ for $k>r/\epsilon$ and hence is not superfast unless $\epsilon\gg n/r$; the algorithm, however, is highly competitive in practice, particularly for SPSD kernel matrices, unlike that of [38] (see [2, 11, 65] and the references therein for kernel matrices and their applications).
@@
The sublinear cost bounds of [7, 38, 11] do not cover superlinear cost of a posteriori estimation of their output error norms and correctness verification. The superfast deterministic algorithm of [33, Part III], however, computes CUR LRA of an $n\times n$ SPSD matrix with both spectral and Frobenius error norms within a factor of $n$ from optimal and as by-product, at no additional cost, estimates the output error norm, verifying correctness.
Chiu and Demanet in [9] defined CUR generator by sampling rows and columns of an input matrix $M$ uniformly at random and proved that this algorithm, which we call Primitive, is quite accurate under the Frobenius norm whp provided that it is directed by a proper “regularization parameter” $\delta$ and that there exists an accurate LRA $XYZ\approx M$, e.g., defined by the SVD of $M$, such that the matrices $X$ and $Z^{*}$ have orthonormal columns and are incoherent.
The algorithm is superfast except for the computation of proper regularization parameter $\delta$, and a numerical example in [9, Sec. 4.1] shows that without this regularization, the approximation error can blow up. Likewise, in our tests in Sec. 6 the output error norm of that algorithm performed without regularization exceeded the norms of all other tested CUR LRA algorithms by factors ranging from 100 to 5,000 in the case of $n\times n$ inputs for $n=256,512,1024$.
In [9, Sec. 3.2] superfast Interpolative Decomposition of [30, Sec. 3.2.3] based on QR factorization with column pivoting is added to that algorithm, and then the authors proved that their estimates for the output accuracy still hold where just one of the factors $X$ or $Z^{*}$ has orthonormal columns and is incoherent.
Cortinovis and Ying in [15] make a little weaker assumptions about the matrices $X$ and $Z^{*}$ – they allow some of their column vectors to be sparse rather than incoherent provided that CUR LRA algorithm incorporates another variant of the Interpolative Decomposition, used earlier in [34, 65]; we outline it in the next subsection. CUR LRA algorithm of [15] was quite accurate in the tests in [15] but, as in [9], is not superfast – it relies on non-canonical transition from CUR generator to CUR LRA based on (2.2).
@@
Although the algorithms of [9, 15] are not superfast, analysis of their output accuracy is still interesting, but the provision of incoherence of the factors of an unknown LRA of $M$ does not hold for a large class of inputs and generally cannot be verified superfast for a fixed $M$.
Our Main Results 1 and 2 (Thms. 4.1 and 9.1) ensure weaker output accuracy than [9, 15] but are universal – they apply to nearly all input matrices and either to all CUR LRAs or to all sketching LRA algorithms, respectively.
1.4.2. Heuristic superfast LRA. Next we recall some superfast LRA algorithms, whose output accuracy was empirically stronger than our Main Results ensure, similarly to what we observed in our tests.
C-A iterations is the most popular example of such algorithms. For two decades of their worldwide application they have been consistently computing accurate CUR LRA for a large class of matrices (see [23, 3, 40, 1], and the bibliography therein).
The heuristic superfast [34, Alg. 2.2] is much less known but important as well. Its output CUR LRA is quite accurate for a large class of unsymmetric $n\times n$ matrices $M$ according to the test results in [34].
For a target rank $r$ this algorithm seeks column and row sets that define the factors $C$ and $R$ of CUR. Initially the two sets are empty and then are recursively increased “a few times” based on nontrivial extension of the algorithm of [9, Sec. 3].
Namely, one first uniformly samples $q>r$ rows of $M$ to obtain a $q\times n$ submatrix, then applies to this submatrix QR factorization with pivoting as in [9], and finally appends the first $r$ column indexes defined by this factorization to the current set of column indexes.
One expands the current set of row indexes by applying the same algorithm to the transpose $M^{*}$.
When this recursive computation of two factors $C$ and $R$ of the CUR of $M$ is completed, a factor $U$ is computed by means of solving a generalized LLSP defined by $M$, $C$ and $R$. This heuristic algorithm involves order of $n^{3/2}$ memory cells and flops, and the authors specify a direction to decreasing that bound to order of $n\log(n)$ but admit: “We have not been able to quantify the error and success probability rigorously for this procedure at this point.”
Instead of QR factorization with column pivoting, Jianlin Xia in [65] uses SRR QR factorization of [19] and calls this technique “progressive alternating direction pivoting”. He combines it with his additional techniques, in particular, randomized error estimation. His algorithms, “numerically shown to have nearly linear complexity” in $m+n$, have consistently output accurate LRAs of real world $m\times n$ kernel matrices $M$ in his extensive numerical tests. He elaborates upon various techniques used for devising his algorithms, but not as much upon their formal support and estimates for their output accuracy and complexity. He admits: “Although a fully rigorous justification of the accuracy is lacking, we give different perspectives to motivate and support the ideas”, but it is still a challenge to specify formally the large subclass of kernel matrices for which his algorithms output accurate LRA. He cites [33, 49] as his close predecessors, but neither his algorithms nor [34, Alg. 2.2] significantly overlap with our current ones for CUR LRA and involve no random sketching at all.
1.4.3. Our old preprints. In Parts I and II we extend the unpublished preprints [45, 46, 47, 48], which cite [50, 51, 52] as their predecessors, where matrix computations under the dual probabilistic model have been formally studied, apparently for the first time. [^8] Such a model was implicit in classical initialization of the Power method with multiplication by a coordinate vector but has not been formally studied or even introduced. The cited preprints stated motivation for further study of superfast LRA as their major goal (which we share) and seemed to be too unorthodox when they appeared, but lately, encouraged by the success of [65, 15], we revisited preprints [47, 48] and extracted from them our Parts I and II, respectively. To present our work in proper historical context, we keep the results of the numerical tests of 2019 intact and do not extend them, e.g., to the numerically stable algorithm of [39], even though this could have strengthened their performance.